OpenStudy (anonymous):

65) The measure of one angle of a regular polygon is 135 degrees. Find the nuber of sides. a) 4 b) 6 c) 8 d) 9

Parth (parthkohli):

See, if the interior angle is 135, then the exterior is 180 - 135 = 45. We know that the exterior angles in a polygon add up to 180. So, number of sides = \(\Large \color{purple}{\rightarrow {180 \over 45} = 4 }\)

OpenStudy (mertsj):

Perhaps the sum of the exterior angles of a polygon add up to 360 since we have to go all the way around the circle.

Parth (parthkohli):

360 I meant :P I got it wrong

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow {360 \over 45} = 8 }\)

Parth (parthkohli):

8 sides actually

OpenStudy (mertsj):

\[135=\frac{(n-2)(180)}{n}\] \[135n=180n-360\] \[-45n=-360\] \[n=8\]